40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1805     Entry time: Wed Jul 29 12:14:40 2009
Author: pete 
Type: Update 
Category: Computers 
Subject: RCG work 

Koji, Pete 

Yesterday, Jay brought over the IO box for megatron, and got it working.  We plan to firewall megatron this afternoon, with the help of Jay and Alex, so we can set up GDS there and play without worrying about breaking things.  In the meantime, we went to Wilson House to get some breakout boards so we can take transfer functions with the 785, for an ETMX controller.  We put in a sine wave, and all looks good on the auto-generated epics screens, with an "empty" system (no filters on). Next we'll load in filters and take transfer functions.

Unfortunately we promised to return the breakout boards by 1pm today.  This is because, according to denizens of Wilson House, Osamu "borrowed" all their breakout boards and these were the last two!  If we can't locate Osamu's cache, they expect to have more in a day or two.

Here is the transfer function of the through filter working at 16KHz sampling. It looks fine except for the fact that the dc gain is ~0.8. Koji is going to characterize the digital down sampling filter in order to try to compare with the generated code and the filter coefficients.


Attachment 1: TF090729_1.png  60 kB  Uploaded Wed Jul 29 13:39:07 2009  | Hide | Hide all
TF090729_1.png
Attachment 2: TF090729_1.png  60 kB  Uploaded Wed Jul 29 13:40:07 2009  | Hide | Hide all
TF090729_1.png
ELOG V3.1.3-